OpenStudy (anonymous):

4cos(x+(pi/3))= A) 2sqrt(3)cosx - 2sinx B) 2cosx - 2sqrt(3)sinx C) 2cosx + 2sqrt(3)sinx D) 2sqrt(3)cosx + 2sinx E) 4cos(x)+2

OpenStudy (raden):

use the formula : cos(A+B) = cos(a)cos(B) - sin(A)sin(B) to get the value of cos(x+(pi/3))

OpenStudy (raden):

then multiply by 4

OpenStudy (anonymous):

so cos(x)cos(pi/3) = (1/2)cosx-(sqrt3/2)sinx times 4 = B?

OpenStudy (raden):

yes, the answer is B

OpenStudy (anonymous):

ok thanks